# Multi-view Graph Structural Representation Learning via Graph Coarsening

### v1

@@ 5.4 Hyper-parameters Study
In this section, we explore the hyper-parameters of our model. We conduct experiments with different settings for 4 hyper-parameters: $K$, $T$, $N$, and hidden size on 3 datasets: IB, IM, and PTC. Figure 4 shows the results of hyper-parameters studies.
[…]
(a) $K$: hierarchy depth
[…]
(b) $T$: the number of layers
[…]
(c) $N$: sampling neighbors
777788889999101010100.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9Hidden Size ($2^{x}$)AccuracyIBIMPTC
(d) Hidden Size
Figure 4: Hyper-parameters studies. Best viewed in color.
Hierarchy Depth $K$ and Layers $T$.
@@
As shown in Figure 4a and Figure 4b, MSLgo has different choices in the hierarchy depth $K$ and the number of Transformer Encoder layers $T$. In short, a deeper and more complicated model improves model performances and enhances the ability to capture complex structural information. However, as the hierarchy extends continuously, overfitting leads to a decline in classification accuracy.
Sampling Neighbors $N$ and Hidden Size.
Figure 4c and Figure 4d show the performance of MSLgo when sampling different numbers of neighbors and selecting different hidden layer sizes, respectively. We can see from the results that the larger the number of sampling neighbors and hidden size, the better the model performance. It indicates that the increase in the sampling and hidden size within a particular range can promote the model to learn high-level structures thoroughly and clearly.
5.5 Case Study
Refer to captionRefer to caption
(a) Case of IB dataset
Refer to captionRefer to caption
(b) Case of N1 dataset
Figure 5: Case studies on IB and N1 datasets. Three graphs are the origin, the coarsening, and the conversion, from left to right. The heatmap below illustrates the weight matrix of the classifier, where the x-axis denotes the feature dimensions, and the y-axis corresponds to three graphs from 0 to 5 (each takes up two rows). Best viewed in color.
@@ 5.5 Case Study
We conduct case studies on IB and N1 datasets to cover each category and explore the choices of the classifier after MSLgo. The results are in Figure 5.
As shown in Figure 5a, high-weighted dimensions gather in the coarsening graph view, contributing more scores to the prediction. It reveals the case of prominent clique structures in social networks like IB. Meanwhile, we find the centrality of the conversion graph view in Figure 5b, which states the fitness of line graphs with edge-central modeling when meeting loops and chains in bioinformatics like N1.
@@
In addition to unique structures, we can still find in the heatmap that the classifier retains weights for some feature dimensions in the original graph view. It is because, after the constrained coarsening and conversion, some nodes retained in the graph still hold valuable structural information, such as the bridge node connecting two structures and the edge node connecting inside and outside a clique. They should be paid equal attention to those unique components.
6 Conclusion
@@
In this paper, we propose a novel multi-view graph structural representation learning model via graph coarsening (MSLgo) on GT architectures for graph classification task. We focus on loops and cliques and investigate the feasibility of treating particular structures node-like as a whole. We build three unique views via graph coarsening and line graph conversion, which helps to learn high-level structural information and strengthen relative position information. We evaluate the performance of MSLgo on six real-world datasets, and the experimental results demonstrate that MSLgo can outperform SOTAs in graph classification task.

### v2

@@ 5.5. Hyper-parameters Study
In this section, we explore the hyper-parameters of our model. We conduct experiments with different settings for 4 hyper-parameters: $K$, $T$, $N$, and hidden size on all eight datasets. Table 6 shows the optimal parameter settings, and Figure 8 shows the results of hyper-parameters studies. Generally, the optimal parameter settings are not the same for different datasets due to various data characteristics. For a single hyper-parameter, we can observe some patterns in results.
@@
As for $K$ and $T$, a deeper and more complicated model improves model performances and enhances the ability to capture complex structural information. However, as the hierarchy extends continuously, overfitting leads to a decline in classification accuracy. The most striking example is the plot of CO and DD on $T$, where the performance of GRLsc decreases substantially when the number of layers changes from 3 to 4.
As for $N$ and $|H|$, in most cases, the larger the number of sampling neighbors and hidden size, the better the model performance. It indicates that the increase in the sampling and hidden size within a particular range can promote the model to learn high-level structures thoroughly and clearly. Among all datasets, IM, PRO, and N109 do not conform to the rules summarized above. IM and PRO can capture long-distance structural information with the help of deeper hierarchy depth without sampling more neighbors. N109 achieves a higher neighbor coverage with only a small amount of sampling for its multi-loop distribution (see Section 5.6 below for more details).
5.6. Case Study
(a) CO (5)
(b) DD (0)
(c) IB (5)
(d) IM (268)
(e) N1 (45)
(f) N109 (540)
(g) PTC (7)
(h) PRO (6)
Figure 9. Case study for all eight datasets. Each case contains three graphs from left to right: the original graph view, the graph coarsening view, and the line graph conversion view. The heat plot below them is the classifier weights, with the x-axis representing the feature dimension and the y-axis representing the three graphs (each graph view takes rows of the number of graph classes).
@@ 5.6. Case Study
We conduct case studies on all eight datasets to fully cover each category and explore the choices of the classifier after GRLsc. The results are in Figure 9.
It can be seen from the cases intuitively that different datasets have different structural characteristics. Even datasets in the same category also have differences in representation. For example, (a) CO, (c) IB, and (d) IM have strong clique structure distribution, graphs in (h) PRO contain multiple connected branches and the similar datasets (e) N1 and (f) N109 consist of multi-chain and multi-loop structure, respectively, and so on. GRLsc focuses on loops and cliques and thus performs well on datasets rich in such structures. As for loops, we take (f) N109 as an example, which contains three benzene rings connected in turn. LCC can identify such loop structures and coarsen them into three supernodes. The remaining two independent carbon atoms are then converted by LGC into a triangle, as shown in the figure. We can see from the heatmap that both LCC and LGC contain vital dimensions contributing most to the classification weights, where LCC in (35, 3), LGC in (6, 4), (24, 5), etc.
As for cliques, we take (c) IB as an example. It forms 12 independent cliques centered on the actor or actress represented by the purple node. A clique represents a collection of actors or actresses from one scene, and some actors or actresses may appear in more than one scene. LCC weakens the concept of nodes but highlights the clique structure by coarsening, strengthening the connection between scenes with corporate actors or actresses. LGC further emphasizes the structural information of the original graph. A clique under the edge-centric view shows a new representation of the purple center node. Compared to the previous case, for datasets with prominent clique structures, LCC contributes more to the weight than LGC, where LCC in (2, 2), (21, 3), (63, 3), and LGC in (13, 5), (46, 5), and so on.
In addition to examples with remarkable features, the dataset also contains (b) DD with a more complex structure and (e) N1 without loops and cliques. We can observe the unique role of LCC and LGC from them. The former coarsens the loops and cliques via LCC and significantly reduces the graph scale, while the latter shortens long chains and marks the position of special forks, such as claw-like structures, through LGC.
Across all eight datasets, we notice that (h) PRO has more unique properties. The original graph is possibly not connected, containing some independent connected branches or components. LCC and LGC start from one particular node, traversing the node and edge set completely. Therefore, multiple connected branches do not affect the operation of our algorithm.
@@
Finally, we can still find in the heatmap that the classifier retains weights for some feature dimensions in the original graph view. It is because, after the constrained coarsening and conversion, some nodes retained in the graph still hold valuable structural information, such as the bridge node connecting two structures and the edge node connecting inside and outside a clique. They should be paid equal attention to those unique components.
6. Conclusion and Future Work
6.1. Conclusion
@@
In this paper, we propose a novel multi-view graph representation learning model via structure-aware searching and coarsening (GRLsc) on GT architectures for graph classification task. We focus on loops and cliques and investigate the feasibility of treating particular structures node-like as a whole. We build three unique views via graph coarsening and line graph conversion, which helps to learn high-level structural information and strengthen relative position information. We evaluate the performance of GRLsc on six real-world datasets, and the experimental results demonstrate that GRLsc can outperform SOTAs in graph classification task.
6.2. Future Work
Though GRLsc achieves remarkable results, we still have a long way to go. In the future, there are two main directions for discussion. First, GRLsc prefers to deal with static graphs at present, while graphs in the real world are constantly changing over time. We try to introduce dynamic graphs. Second, GRLsc focuses on two specific structures: loops and cliques. However, graph structures are more complex and diverse. We will consider extending to general structures to mine richer information at a high level.